OpenStudy (anonymous):

Evaluate if x =3 and y = -2: x2 + 2y A. 5 B. 10 C. 2

OpenStudy (anonymous):

Does x2 mean

OpenStudy (anonymous):

x^2 + 2y or 3^2 + 2(-2) or 9 - 4 = 5

OpenStudy (anonymous):

3^2=9 -2*2=-4 9-4=5 (a)